# DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models

### v2

@@
Abstract
Multimodal large language models (LLMs) have demonstrated significant potential in a wide range of AI applications. Yet, training multimodal LLMs suffers from low efficiency and scalability, due to the inherent model heterogeneity and data heterogeneity across different modalities.
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal large language models on large-scale clusters. The core of DistTrain is the disaggregated training technique that exploits the characteristics of multimodal LLM training to achieve high efficiency and scalability. Specifically, it leverages disaggregated model orchestration and disaggregated data reordering to address model and data heterogeneity respectively. We also tailor system optimization for multimodal LLM training to overlap GPU communication and computation. We evaluate DistTrain across different sizes of multimodal LLMs on a large-scale production cluster with thousands of GPUs. The experimental results show that DistTrain achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput. The ablation study shows the main techniques of DistTrain are both effective and lightweight.
1 Introduction
Recent advances in large language models (LLMs) are catalyzing a new wave of AI applications [1]. However, LLMs are predominantly text-based, restricting their ability to understand and generate multimodal content. Emerging multimodal LLMs address this gap by integrating various modalities such as texts, images, and audios into LLMs which significantly enhances LLM’s applicability. Multimodal LLMs demonstrate great potential in tasks like image understanding [2, 3, 4, 5], audio comprehension [6, 7], and embodied AI [8, 9]. Many organizations are actively developing their multimodal LLMs, such as OpenAI’s GPT-4o [10], Google’s Gemini [11] and PaLM-E [9], Meta’s Chameleon [12], etc.
Training multimodal LLMs demands vast computational resources. According to the scaling law [13], model size and training data volume are crucial for determining model capabilities. Substantial efforts are invested in training models with billions of parameters on trillion-scale tokens. For instance, Meta’s Chameleon [12] is a 34B multimodal LLM trained on more than 4.8 trillion tokens. Leading organizations often deploy large-scale clusters, equipped with thousands of GPUs, for such training tasks. One technical report [14] reveals that training a GPT-3 175B with 300 billion tokens costs $4.6 million and lasts one year using 355 V100 GPUs. Consequently, it is crucial to develop an efficient and scalable training framework to minimize costs and accelerate training.
@@ 1 Introduction
Figure 1 depicts the mainstream model architecture of multimodal LLMs, comprising three primary modules: the modality encoder, LLM backbone, and the modality generator [15, 16]. These modules are linked by the projector, which may incorporate MLP or cross-attention layers. The modality encoder transforms input from various modalities into a unified embedding space. The embedding vectors are then analyzed by the LLM backbone, a transformer model, to discern data patterns and inter-modal relationships. Subsequently, the modality generator translates this processed information back into coherent outputs tailored to each modality.
Existing LLM training framework, e.g., Megatron-LM [17], can be extended to train multimodal LLMs by treating the multimodal modules as additional layers within the LLM. However, training multimodal LLMs poses two substantial challenges: model heterogeneity and data heterogeneity. The fundamental issue of model heterogeneity stems from the need to process diverse modalities with different modules that vary dramatically in size and operator complexity. The model heterogeneity across different modules (i.e., modality encoder, LLM backbone, and modality generator) introduces severe pipeline bubbles, resulting in poor GPU utilization. Meanwhile, data heterogeneity bursts onto the scene due to the intricate and unstructured nature of multimodal input data. The data heterogeneity across different modality data leads to inter-microbatch and intra-microbatch training stragglers, which prolong the training duration and exacerbate the pipeline bubbles. These challenges collectively limit the efficiency and scalability of multimodal LLM training, resulting in MFU as low as ~20% in production-level training (§8.1).
To this end, we present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs. DistTrain achieves state-of-the-art MFU which is close to unimodal LLM training and effectively scales to large clusters with thousands of GPUs. The core principle of DistTrain is disaggregated training, including GPU training disaggregation and CPU preprocessing disaggregation, which facilitate to address model and data heterogeneity respectively.
@@
For model heterogeneity, we meticulously analyze the pipeline bubbles stemming from model heterogeneity and identify their root causes (§2.2). The GPU training disaggregation of modality encoder, LLM backbone, and modality generator enables adaptive orchestration across the three modules in multimodal LLM. Building on this, we propose disaggregated model orchestration that navigates the complicated design space to choose the optimal resource and parallelism configurations. This innovative approach minimizes the pipeline bubbles caused by model heterogeneity and achieves optimal training efficiency.
For data heterogeneity, we categorize the training stragglers into inter-microbatch and intra-microbatch stragglers (§2.3). The CPU preprocessing disaggregation allows efficient data preprocessing (e.g., data decompression and reordering) with negligible runtime overhead. From this foundation, we incorporate disaggregated data reordering into the preprocessing to strategically reorder training data without additional overhead. The inter-microbatch reordering algorithm reorders the data samples to evenly distribute the load across different data parallelism groups. The intra-microbatch reordering algorithm reorders the microbatches tailored for (interleaved) 1F1B pipeline scheme to minimize the pipeline bubbles. This two-level approach effectively mitigates data heterogeneity.
In addition, we customize system optimization for multimodal LLM training. We implement an in-house collective communication library, StepCCL, to hide the communication overhead within the computation.
@@
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity. It delivers state-of-the-art MFU on large-scale clusters with thousands of GPUs.
We identify and discuss the primary challenges associated with multimodal LLM training, which are summarized as model heterogeneity and data heterogeneity.
@@
We propose disaggregated training for multimodal LLM. It leverages disaggregated model orchestration and disaggregated data reordering to effectively address model and data heterogeneity, respectively.
@@
We implement DistTrain and conduct experiments on our production cluster with thousands of GPUs. The experimental results show that DistTrain achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
2 Motivation
2.1 Multimodal Large Language Model Training
Large language model. Large language models (LLMs) [18, 19, 20] have revolutionized natural language processing (NLP) by achieving state-of-the-art performance on a wide range of tasks, such as text generation, translation, and summarization. Many organizations have raced to develop their own LLMs, such as OpenAI’s GPT-4 [21], Google’s Gemini [11], and Meta’s Llama [22]. The core architecture of LLMs consists of a stack of homogeneous transformer layers [18] that use self-attention mechanisms to capture contextual information in text. LLMs are pre-trained with unsupervised learning on large-scale text corpora and then fine-tuned on task-specific text datasets. The text data is typically tokenized into fixed-length sequences, which are then fed into the model to learn the underlying patterns. The training process involves weeks or even months of computation on dedicated AI clusters with thousands of GPUs. According to one technical report [14], training 175B GPT-3 requires 4.6 million dollars. Optimizing the training process is essential to reduce the stupendous cost and accelerate the model deployment.
Multimodal LLM. The unimodal LLMs are limited to processing text data, which restricts their applicability to multimodal tasks (e.g., image understanding and generation). As a result, multimodal LLMs have emerged to address this limitation by integrating multiple modalities (e.g., images, audios, and videos) into the advanced LLMs [15], which support multimodal inputs and outputs during LLM generation. For example, GPT-4o [10] garners widespread attention by facilitating more natural interactions with humans through both visual and auditory modalities. Moreover, the predominantly text-based data in human societies is finite [23]. Harnessing multimodal data is inevitable to continually expand and enhance LLM capabilities.
@@ 2.1 Multimodal Large Language Model Training
Figure 1 illustrates the model architecture of a multimodal LLM, which consists of three modules: a modality encoder, an LLM backbone, and a modality generator [24, 15, 16]. The modality encoder transforms input data from different modalities (e.g., ViT [25] for images and Beats [26] for audios) into an intermediate representation (i.e., an embedding tensor), which is then projected into a unified embedding space across modalities with input projection layers (e.g., MLP and cross-attention). The LLM backbone, typically a transformer model (e.g., GPT [19, 20] and Llama [22]), processes the multimodal embeddings to discern the intricate data patterns and inter-modal relationships. The output data of the LLM backbone is subsequently refined by output projection layers, which tailor the information for each modality. Finally, the modality generator (e.g., Diffusion [27] for images and AudioLDM [28] for audios) transforms the LLM-processed information back into the respective modal outputs for generation.
@@
Multimodal LLM training necessitates training all three modules simultaneously. Additionally, during different training phases, specific modules are frozen to stabilize training loss and enhance model effectiveness [15]. Regarding the training data, the input sequence comprises text, image, and audio tokens. These data from different modalities are tokenized into subsequences which are then interleaved to form fixed-length training sequences [12]. In cases where the input sequence falls short, it is padded with zero tokens for batching. Distinct modality sub-sequences are demarcated by special tokens and processed through modality-specific encoders.
Training framework. To train the multimodal large language model with large-scale clusters, the de facto solution is to leverage Megatron-LM, a highly efficient and robust training framework for large-scale transformer models. Megatron-LM employs a unified parallelism strategy for the entire model. It combines tensor parallelism (TP) and pipeline parallelism (PP) to distribute the model parameters across multiple GPUs. TP divides the model parameter within each layer, while PP partitions parameters between layers. It also leverages data parallelism (DP) to distribute the training data. For multimodal LLMs, Megatron-LM is easily extended to integrate additional multimodal modules. Figure 2 illustrates the training topology of multimodal LLMs with Megatron-LM. Specifically, Megatron-LM treats the multimodal modules as additional layers within the LLM and incorporates additional PP stages to accommodate the modality encoder and generator. The same TP strategy used in the LLM backbone is applied to these two multimodal modules. If the modality encoder and generator are not large enough, they are replicated across the GPUs in the TP group to maximize resource utilization. As for DP, Megatron-LM applies the same DP strategy to the multimodal modules as the LLM backbone. The projector is co-located with the modality encoder and generator and is replicated across the GPUs in the TP group.
However, this training framework introduces significant computation imbalance stemming from model heterogeneity and data heterogeneity due to its rigid model orchestration method. (i.e., the multimodal modules share the same DP and TP strategies with LLM backbone). It only achieves ~20% MFU (§8.1) which is significantly lower than the MFU of ~50% observed in training unimodal (text-only) LLMs [29].
2.2 Model Heterogeneity
@@
The first challenge is the computation imbalance arising from model heterogeneity. Each module in multimodal LLMs bears different computational demands due to varying operators and inputs. For instance, ViT, as modality encoder, is constructed with narrow transformer layers (i.e., small hidden size), whereas LLM backbone is built on broader transformer layers (i.e., large hidden size). Meanwhile, Diffusion, as modality generator, utilizes a combination of convolution and attention layers (i.e., U-Net). This architectural diversity results in distinct computation time for each module. Figure 3 shows varying forward time under different input configurations with Megatron-LM. We demonstrate one PP stage of LLM backbone with PP size of 10 and TP size of 8. The first configuration parameter is the number of images in the 8K input sequence, and the second is the image resolution. The time differs markedly across different configurations.
@@
The computational imbalance between modules leads to two types of pipeline bubbles in pipeline parallelism. The first arises in the modality encoder and generator stages, as shown in Figure 4(a), resulting from their inadequate utilization of assigned GPU resources. The second type emerges in the stages of the LLM backbone, as shown in Figure 4(b). This is because the intensive computational demands of the encoder and generator extend their stage durations. Due to the pipeline dependency, the LLM stages are forced to wait for the multimodal stage to complete, thereby creating pipeline bubbles. The latter problem is particularly pronounced during large-scale multimodal LLM training, where the bulk of GPU resources are allocated to the LLM backbone. These pipeline bubbles, stemming from model heterogeneity, substantially diminish the MFU during the training.
2.3 Data Heterogeneity

### v3

@@
Abstract.
Multimodal large language models (LLMs) empower LLMs to ingest inputs and generate outputs in multiple forms, such as text, image, and audio. However, the integration of multiple modalities introduces heterogeneity in both the model and training data, creating unique systems challenges.
We propose DistTrain, a disaggregated training system for multimodal LLMs. DistTrain incorporates two novel disaggregation techniques to address model and data heterogeneity, respectively. The first is disaggregated model orchestration, which separates the training for modality encoder, LLM backbone, and modality generator. This allows the three components to adaptively and independently orchestrate their resources and parallelism configurations. The second is disaggregated data preprocessing, which decouples data preprocessing from training. This eliminates resource contention between preprocessing and training, and enables efficient data reordering to mitigate stragglers within and between microbatches caused by data heterogeneity. We evaluate DistTrain across different sizes of multimodal LLMs on a large-scale production cluster. The experimental results show that DistTrain achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on training throughput.
ACM Reference Format:
Zili Zhang, Yinmin Zhong, Yimin Jiang, Hanpeng Hu, Jianjian Sun, Zheng Ge, Yibo Zhu, Daxin Jiang, Xin Jin. 2025. DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models. In SIGCOMM ’25, August 8-August 11, 2025, Coimbra, Portugal, 15 pages.
DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training …
1. Introduction
Traditional multimodal models, which focus solely on either cross-modal feature representation (e.g., CLIP (Radford et al., 2021)) or multimodal content generation (e.g., Stable-Diffusion (Rombach et al., 2022) and VAE (Kingma et al., 2013)), cannot simultaneously process multimodal inputs and generate multimodal outputs. This restricts their capabilities and hinders seamless human interaction. Multimodal large language models (LLMs) emerge to integrate the strengths of LLMs (Vaswani et al., 2017; Brown et al., 2020; Touvron et al., 2023) with traditional multimodal models to simultaneously ingest multimodal inputs and generate multimodal outputs in auto-regressive manner. Multimodal LLMs have already shown great potential in tasks such as vision understanding and generation (Liu et al., 2024b; Zhang et al., 2023b; Wang et al., 2024; Dong et al., 2023; Kondratyuk et al., 2023), audio comprehension (Rubenstein et al., 2023; Borsos et al., 2023), and embodied AI (Zhang et al., 2023a; Driess et al., 2023). Many organizations are actively developing multimodal LLMs, such as OpenAI’s GPT-4o (gpt, 2024), Google’s Gemini (gem, 2024), and Meta’s Llama3 (lla, 2024), etc.
@@ 1. Introduction
Figure 1 depicts the model architecture of multimodal LLMs, comprising three modules: modality encoder, LLM backbone, and modality generator (Dong et al., 2023; Yin et al., 2023; Zhang et al., 2024). We emphasize this architecture underpins most state-of-the-art multimodal LLMs (Table 1). These modules are linked by projectors, which may incorporate MLP or cross-attention. The modality encoder transforms inputs from various modalities into different modality tokens. These tokens from different modality are interleaved as a sequence for LLM training. The modality generator translates the processed information back into coherent outputs tailored to each modality.
Training multimodal LLMs is more challenging than training traditional multimodal models. Traditional multimodal model training frameworks like DistMM (Huang et al., 2024) lack support for simultaneous training of encoders, LLMs, and generators, particularly techniques tailored for LLMs. The current practice for training multimodal LLMs (Alayrac et al., 2022; Wu et al., 2024) extends LLM training frameworks (e.g., Megatron-LM (Shoeybi et al., 2019)) with additional modality encoder and generator modules, which uses a monolithic approach. In this context, the monolithic training refers to two aspects. First, it uses the same data and tensor parallelism strategies of distributed training across the modality encoder, LLM backbone, and modality generator modules. As for pipeline parallelism, the encoder and generator modules are incorporated as additional pipeline stages in the training pipeline. Second, it couples data preprocessing into training process, and co-locates data preprocessing and training on the same machines.
The monolithic approach falls short of handling unique systems challenges in multimodal LLM training. Compared to traditional LLM training that only handles text, the multiple modalities in multimodal LLM training introduces model heterogeneity and data heterogeneity. Model heterogeneity stems from processing multiple modalities with different modules (i.e., encoder, LLM backbone, and generator) that vary dramatically in size and operator complexity. Using the same parallelism strategies across different modules introduces severe pipeline bubbles, resulting in poor GPU utilization (§2.2). Data heterogeneity emerges from processing multimodal data that differ significantly in volume and processing cost, e.g., a high-resolution image is much larger and requires more resources for preprocessing than a line of text. In the monolithic approach, where data preprocessing and training are integrated and co-located, the substantial processing cost significantly hinders the training process. Additionally, the heterogeneity across different modality data (i.e., each training sample has a varying number of modality tokens) leads to inter-microbatch and intra-microbatch training stragglers, which prolong the training duration and exacerbate the pipeline bubbles (§2.3). In our initial efforts using the monolithic approach, we observed extremely low MFU in production training, as low as 20% (§7.1).
To this end, we propose DistTrain, a disaggregated training system for multimodal LLMs. DistTrain applies the systems principle of disaggregation in the context of multimodal LLM training to address the unique systems challenges. We design two novel disaggregation techniques, which are disaggregated model orchestration and disaggregated data preprocessing, to address model and data heterogeneity, respectively.
@@
For model heterogeneity, we analyze the pipeline bubbles caused by multiple modalities across different modules in multimodal LLMs, and identify their root causes (§2.2). To address this, we design disaggregated model orchestration to separate the training of modality encoder, LLM backbone, and modality generator. This separation enables adaptive orchestration of resources and parallelism configurations across the three modules. Building on this, we formulate the model orchestration as an optimization problem aimed at minimizing the training time per iteration. We then design an adaptive model orchestration algorithm that navigates the complicated search space to find the optimal resource and parallelism configurations. Disaggregated model orchestration is able to minimize pipeline bubbles caused by model heterogeneity and achieves optimal training efficiency.
For data heterogeneity, we characterize the training data of different modalities, and unveil their preprocessing overheads and the stragglers they brought into the training process (§2.3). We categorize the stragglers into intra-microbatch and inter-microbatch stragglers. We design disaggregated data preprocessing to eliminate the interference between preprocessing and training. Importantly, such disaggregation enables flexible data preprocessing, and provides new opportunities for data reordering to mitigate stragglers without additional overhead. We use intra-microbatch reordering to evenly distribute the load across data parallelism groups, and inter-microbatch reordering to hide the training stragglers into the training pipeline. This dual-level reordering effectively mitigates the stragglers caused by data heterogeneity. Importantly, since the reordering only permutes the sequence of commutative gradient accumulation operations, it preserves the training’s convergence semantics.
@@
We identify the unique system challenges introduced by multiple modalities in multimodal LLM training, and summarize them as model and data heterogeneity.
@@
We propose DistTrain, a disaggregated training system for multimodal LLMs, and design two novel techniques—disaggregated model orchestration and disaggregated data preprocessing—to address model and data heterogeneity.
@@
We implement DistTrain and conduct experiments on our production cluster with thousands of GPUs. The experimental results show that DistTrain achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on training throughput.
2. Motivation
2.1. Multimodal LLM Training
Traditional multimodal models. Traditional multimodal models encompass diverse architectures, such as contrastive learning (Radford et al., 2021) and stable diffusion (Rombach et al., 2022) (SD). Contrastive learning uses encoder models to capture cross-modal similarities, primarily for feature representation. Stable diffusion generates images from text prompts by iteratively refining random noise through a denoising process, enabling text-to-image generation. However, these models are task-specific and lack the flexibility to handle diverse multimodal scenarios, which cannot process multimodal inputs and generate multimodal outputs simultaneously, restricting their model capabilities in complex real-world settings and hindering seamless interaction with humans.
Multimodal | Encoder(s) | LLM | Generator(s) |
LLM | Backbone | ||
Flamingo (Alayrac et al., 2022) | NFNet (Brock et al., 2021) | GPT-3 (Brown et al., 2020) | LM-Head |
LLaVA (Liu et al., 2024a) | CLIP (Radford et al., 2021) | Vicuna (chi, 2023) | LM-Head |
PaLM-E (Driess et al., 2023) | ViT (Dosovitskiy, 2020) | PaLM (Chowdhery et al., 2023) | LM-Head |
EMU (Sun et al., 2023b) | EVA-CLIP (Sun et al., 2023a) | Llama (Touvron et al., 2023) | LM-Head, SD (Rombach et al., 2022) |
Bagel (Deng et al., 2025) | ViT (Sun et al., 2023a) | Qwen2.5 (yan, 2024) | LM-Head, VAE (Kingma et al., 2013) |
VideoPoet (Kondratyuk et al., 2023) | MAGViT (Yu et al., 2023), | GPT (Radford et al., 2018) | MAGViT (Yu et al., 2023), |
SoundStream (Zeghidour et al., 2021) | SoundStream (Zeghidour et al., 2021) |
Multimodal LLMs. Large language models (LLMs)(Achiam et al., 2023; gem, 2024; lla, 2024), built on homogeneous transformer layers and trained on extensive text corpora, excel at understanding and generating high-quality text but are limited to text-only processing. Multimodal LLMs address these limitations by integrating the strengths of LLMs with traditional multimodal models, enabling the simultaneous processing of diverse inputs (e.g., images, audio, video) and generating multimodal outputs in auto-regressive manner. For example, GPT-4o (gpt, 2024) gains wide attention by facilitating more natural interaction with humans through both visual and auditory modalities.
@@ 2.1. Multimodal LLM Training
Figure 1 illustrates the model architecture of a multimodal LLM, which consists of three modules: a modality encoder, an LLM backbone, and a modality generator (Alayrac et al., 2022; Yin et al., 2023; Zhang et al., 2024). Different from traditional multimodal models, this generative model facilitates a more robust and interpretable process for understanding and generating multimodal contents through auto-regressive decoding. Specifically, the modality encoder transforms input data from different modalities (e.g., ViT (Dosovitskiy, 2020) for images and Beats (Chen et al., 2022) for audios) into an intermediate representation, which is then converted into different modality tokens with input projector (e.g., MLP and cross-attention). These tokens from different modalities form a sequence for LLM training. The LLM backbone, typically a transformer model (e.g., GPT (Radford et al., 2019; Brown et al., 2020) and Llama (lla, 2024)), processes the input sequence to extract intricate data patterns and inter-modal relationships. The LLM backbone’s output is refined by an output projector and then passed to a modality generator (e.g., Diffusion (Rombach et al., 2022) for images, AudioLDM (Liu et al., 2023a) for audio), which converts the processed information into the corresponding modal outputs. Table 1 summarizes the three modules of some state-of-the-art multimodal LLMs. We omit text tokenizer (e.g., BPE) and encoder (e.g., text embedding) from the table’s encoder column. LM-Head is responsible for decoding the text tokens
@@
Multimodal LLM training necessitates training all three modules simultaneously. Additionally, during different training phases, specific modules are frozen to stabilize training loss (Chen et al., 2025; Deng et al., 2025). Regarding the data, input sequences comprise text, image, and audio tokens. The data from different modalities are encoded into subsequences which are then interleaved to form fixed-length training sequences (Team, 2024).
Training frameworks. Multimodal model training frameworks such as DistMM (Huang et al., 2024) do not support training encoder, LLM backbone and generator simultaneously. Especially, they lack support training techniques (Chen et al., 2024; Narayanan et al., 2021; Rajbhandari et al., 2020) for LLMs. In contrast, LLM training frameworks like Megatron-LM (Shoeybi et al., 2019) can be extended with modality encoder and generator modules, making them widely used in production multimodal LLM training, e.g., Flamingo (Alayrac et al., 2022), DeepSeek-VL2 (Wu et al., 2024) and VideoPoet (Kondratyuk et al., 2023). Figure 2 shows training multimodal LLMs with Megatron-LM. Megatron-LM adds multimodal modules as additional layers, and incorporates additional pipeline parallelism (PP) stages to accommodate encoder and generator. The same tensor parallelism (TP) strategy used in the LLM backbone is applied to encoder and generator. If the encoder and generator are not large enough, they are replicated across the GPUs in the TP group to maximize resource utilization. As for data parallelism (DP), Megatron-LM applies the same DP strategy to the multimodal modules as the LLM backbone. The projector is co-located with the encoder and generator, and is replicated across the GPUs in the TP group. Moreover, data preprocessing is incorporated into the training process and co-located with training on the same node. This approach is monolithic in terms that the encoder and generator use the same parallelism strategy as the LLM backbone, and that preprocessing and training are integrated and co-located. Such a monolithic approach introduces significant computation imbalance stemming from model and data heterogeneity.
2.2. Model Heterogeneity
Characterization. Each module in multimodal LLMs has different computational demands due to varying operators and inputs. For instance, ViT, as modality encoder, is constructed with narrow transformer layers (i.e., small hidden size), whereas the LLM backbone is built with wide transformer layers (i.e., large hidden size). Meanwhile, Diffusion, as modality generator, utilizes a combination of convolution and attention layers. The diversity in model architecture results in distinct computation time for each module. Figure 3 shows varying forward time under different input configurations with Megatron-LM. We demonstrate one PP stage of LLM backbone with PP size of 10 and TP size of 8. The first configuration parameter is the number of images in the 8K input sequence, and the second is the image resolution. The results reveal a key performance disparity: while the LLM backbone maintains a consistent forward time, the modality encoder and generator exhibit significant variations.
@@ 2.2. Model Heterogeneity
Pipeline imbalance. The computation imbalance between modules leads to two types of bubbles in pipeline parallelism. The first arises in the modality encoder and generator stages, as shown in Figure 4(a), resulting from their inadequate utilization of allocated GPU resources. The second type arises in the stages of the LLM backbone, as shown in Figure 4(b). This is because the intensive computation demands of the encoder and generator increase their stage durations. Due to the pipeline dependency, the LLM stages are forced to wait for the multimodal stage to complete, thereby creating pipeline bubbles. The latter problem is particularly pronounced during large-scale multimodal LLM training, where the bulk of GPU resources are allocated to the LLM backbone. These pipeline bubbles, stemming from model heterogeneity, reduce the MFU significantly during training.
2.3. Data Heterogeneity